An electric field E=Axi^ exists in the space where A=10V/m2​ ..​Take the potential at (10m,20m) to be zero. Find the potential at the origin.​

Grade:12

3 Answers

Vikas TU
7 years ago
E = -dV/dx
dV = – Axdx
Integerating both sides,
V = -Ax^2/2 + c
at (10,20)
V = 0 => 150A = C
C = 1500
 
V at origin,
V  =  1500 V/m^2
 
Sourabh Rai
4 years ago
E = -dV/dx and. E = Ax î
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

E = -dV/dx
dV = – Axdx
Integerating both sides,
V = -Ax^2/2 + c
at (10,20)
V = 0 => 150A = C
C = 1500
V at origin,
V = 1500 V/m^2

Thanks and Regards